The larger of two numbers is 5 more than twice the smaller. If the smaller is subtracted from the larger, the result is 12. Find the numbers using elimination method.
Answer by faceoff57(108)

The larger of two numbers is 5 more than twice the smaller. If the smaller is subtracted from the larger, the result is 12. Find the numbers using elimination method.
L = 5 + 2S
L - S = 12
(5 + 2S) - S = 12
5 + 2S - S = 12
5 + S = 12
S = 7
L - 7 = 12
L = 19
Larger number is 19, Smaller number is 7.